ANSWER THE FOLLOWING QUESTIONS

1: Mr. Meadows Cookie Company makes a variety of chocolate chip cookies in the plant in Albion, Michigan. Based on orders received and forecasts of buying habits, it is assumed that the demand for the next three months is 600, 800 and 450, expressed in thousands of cookies. During a 50-day period when there were 80 workers, the company produced 2.4 million cookies. Assume that the number of workdays in each month is 25. There are currently 70 workers employed, and there is no starting inventory of cookies. a. What is the minimum constant workforce (level strategy) required to meet demand (shortages not allowed) over the next three months? b. Assume that the inventory holding cost is 5 cents per cookie per month, hiring cost is $400 per worker, and firing cost is $600 per worker. Evaluate the cost of the plan derived in a.

QUESTION 2: Historical demand for a product is: Month Demand 1 January 40 2 February 43 3 March 45 4 April 48 5 May 50 a. Using a simple three-month moving average, find the April and May forecast. Compute MAD. b. Using a single exponential smoothing with and a May forecast = 49, find the June forecast c. Using a double exponential smoothing method with and, find and. d. Using and found in part, find the June forecast made in January.

QUESTION 3: An analyst predicts that a 70 percent experience curve should be an accurate predictor of the cost of producing a new product. Suppose that the cost of the 2nd unit is $4,200. Estimate the cost of producing a. (3 points) the 4th unit. b. (3 points) the 1st unit. c. (4 points) the 3rd unit.

STATE WHETHER TRUE OF FALSE

21) Depreciation is a source of funds.
22) Working Capital is the ratio of assets to liabilities.
23) Society imposes regulations on businesses to protect individuals from each other.
24) Double-entry accounting entails having one set of books for internal tracking and
another set of books for reporting to stockholders and for paying taxes.
25) The purpose of an income statement is to show whether the company ahs made a
profit or a loss over a defined period.
26) The opposite of a writedown is a gain on the sale of assets.
27) If a company has a product with a margin of 74% and the plant is running at 90%
capacity, a discount of 25% on marginal sales may be acceptable.
28) A short-term lender may require a minimum level of working capital as a condition
that the borrower must meet in order to maintain the credit line.
29) A company purchases a grader for $30,000, and no other assets in that asset class
in that year. The UCC remaining from the previous year is $10,000. For CRA rate of
30%, this year’s CCA is $7,500.

30) The change in retained earnings in a reporting period is the net income minus
dividends.
Answer #1

*****Dear Student, you are violating the Chegg Guidelines, by posting multiple questions in a single post. Please make a note that Revised Chegg Guidelines seriously restrict us in answering only one question! However, I've answered first question so that you would rate this answer positively. DO NOT VIOLATE Chegg guidelines as they may have a huge impact on your account in the future. So please understand and rate this answer with thumbs up kindly. It is a gesture that shows you support me. Thank you******

Productivity per worker for 46 days = (2.4 x 106)/80 = 30,000‬

Productivity per worker per day = 30,000‬/50 = 600 units/worker/day

k = Productivity (in 1000) per worker per day = 0.6 units/worker/day

a.

Month

Net Production Required

(1000)

Cumulative net production required

#Days per month

Production per worker

(in 1000)

Cumulative Production per worker

# of Workers Needed

A

B

C

D = C*k

E

F = B/E

1

600

600

25

15

15

40

2

800

1400

25

15

30

46.67

3

450

1850

25

15

45

41.11

Maximum of F is 46.67 = 47

So , minimum employee required = 47 employees

b.

Month

Beginning Inventory (in 1000's)

Production per worker

(in 1000’s)

Production by 47 workers (in 1000's)

Demand
(in 1000's)

Ending Inventory
(in 1000's)

B.I

E

P = E*47

A

EI = BI + P - A

1

0.00

15

705

600

105

2

105

30

1410

800

715

3

715

45

2115

450

2380

Total = 105 + 715 + 2380 = 3200

Beginning workforce level = 70

Minimum workforce required = 47

Workforce fired = 23

Total firing cost = 23 * 600 = $13,800

Total inventory cost = 0.05 * E.I = 0.05* 2380 * 1000 = $119,000

Total plan cost = $132,800